### v1

# Moiré exciton polaritons in twisted photonic lattices at room temperature

Abstract
Moiré lattices attract intensive attention in the double graphene/TMD layers and photonic crystals due to the interesting exotic physics within these structures. However, precise measurement of the moiré ground states, excited states and Bloch bands in the twisted photonic lattices is still illusive. In this work we report the strong coupling between the excitons of CsPbBr3 microplates and the photonic modes of the moiré lattice at room temperature. Depending on the coupling strength between the nearest potential sites, we observe staggered moiré polariton ground states, excited states trapped in the potential sites and moiré polariton bands across the twisted photonic lattice. In addition, the phase locking of moiré zero (stable in-phase) states and moiré $\pi$ (metastable antiphase) states with different spatial distributions are measured. Moiré polariton distribution can be tuned in the shape of parallelogram by controlling the depth and width of the potential in one photonic lattice with another one fixed. Our work lays the foundation to study moiré exciton polariton Wigner crystals and Luttinger liquid in twisted photonic lattices at room temperature.
Moiré lattices in the graphene bilayer and TMD monolayer heterostructures offer a tunable platform where the interlayer coupling of electrons and excitons leads to plethora of physics phenomena, for example, the transition from superconductor to Mott insulator caoyang , observation of the quantum Hall effect moire quantum hall effect and nontrivial gauge field moire gauge potential , the appearance of the moiré excitons exiton moire1 ; exiton moire2 ; exiton moire3 , the emergence of the correlated exciton insulators flat band bilayer1 ; flat band bilayer2 ; exciton flat band . On the other hand, the bands resulted from the coupling between the moiré potential sites have been extended into the photonic system. In these systems, the localization and delocalization of light and nonlinear optical solitons can be observed by adjusting one photonic lattice against the second one into a particular angle photonic moire lattice . Importantly, the period of the moiré lattice in the twisted photonic structures can reach the size of several $\mu$ms, thus the bands due to the coupling between the moiré potentials can be probed within several $\mu$m-1 in the normal angle-resolved spectroscopy setup. In this case, both the momentum space and real space imaging can be directly measured. For example, photonic bands in the optical frequency range are probed in a double photonic crystals moire band . Cold atom condensates loaded in twisted photonic lattices have also been realized and the transition from superfluid to Mott insulator in the created moiré lattice is shown atom moire lattice .
Different from TMD moiré heterostructures where excited moiré states can be observed by controlling the electron density, excited moiré states is still challenging to study in twisted photonic lattices, where the moiré potential trap is shallow. Deep moiré potentials allow to study confined mode distribution and excited mode coupling induced Bloch band across the twisted photonic lattices. Importantly, phase locking can be observed in the moiré band. For example, the phase synchronization within a flat band can enhance the performance of nanocavity laser array greatly flat band phase . On the other hand, $\pi$ phase locking at the top of the moiré ground bands and bottom of the moiré excited bands can lead to different mode distribution in the momentum space and real space compared to the ground band, which can offer to study quantum simulation with tunable phase modulation.
Figure 1: Schematics of the microcavity and optical image of the moiré lattice. (a) The structure of the microcavity, which consists of the perovskite microplate onto the bottom DBR and a twisted photonic lattice, which is plotted in the right panel with the relative angle of 40. (b) The optical image of the perovskite microplates within the microcavity, where one can see the moiré pattern, the right panel is the zoomed schematics of the moiré pattern.
Exciton polaritons can provide to investigate different phase locking when it is loaded into a photonic lattice yamamoto . As a quasiparticle due to the strong coupling between the exciton and photon mode microcavity book , in Fabry-Perot microcavity, exciton polaritons can demonstrate similar Bose-Einstein condensation process as the cold bosonic particles at much higher temperature due to the very light effective mass polariton BEC1 ; polariton BEC2 . To observe phase locking of zero and $\pi$ of exciton polaritons in the moiré lattice at room temperature, particular semiconductors with robust excitons and easy integration with photonic structures are needed, among which the perovskite CsPbBr3 microplates are studied intensively recently due to the large exciton binding energy and oscillator strength. Room temperature exciton polariton condensation based on CsPbBr3 microplates has been demonstrated in the microcavities Yao RD ; Xiaokun vortex ; gy pvk ; lhh pvk ; polariton pvk1 ; polariton pvk2 .
In this work we realize to observe strong coupling between the excitons in the perovskite CsPbBr3 microplates and moiré photonic modes in a twisted photonic lattice which is composed of two tilted one dimensional lattice fabricated within the microcavity at room temperature. Similar moiré pattern with two identical 1D photonic lattices has been reported in moire4 . In our work, the moiré exciton polariton zero state and $\pi$ state in this twisted photonic lattice are measured directly. The staggered moiré polaritons with different energies are either confined in the potential trap or in the potential barrier region. By modifying the potential trap depth and width of one photonic lattice, moiré polaritons in the shape of parallelogram deformed with isolated dots, dimers and extended lines are observed which is sensitive against the disorder. Our work offers to investigate room temperature quantum simulation and Wigner crystals based on moiré exciton polaritons.
The microcavity is fabricated by transferring the chemical vapor deposition (CVD) grown perovskite microplates CsPbBr3 onto the bottom DBR. The 1D photoresist lattice with the period of 4 $\mu$m (duty circle: 50:50) and the thickness of 200 nm is formed by using the lithography technique onto the top of perovskite microplates (A layer of SiO2 is deposited onto the perovskite for the protection, the details can be found in the Methods section). The same 1D photoresist lattice with the same duty circle and depth is formed onto the top DBR with particular angle against the bottom DBR. The microcavity is fabricated by pasting the top DBR against the bottom DBR using the adhesive with an air gap of around 3 $\mu$m. The schematic graph of the microcavity and the photonic lattice with the twisted angle of around 40 are shown in Figure 1(a), where the moiré stripes consisting of periodic rhombus arrays along horizontal and vertical direction are clearly seen (the right panel of Figure 1(a)). The optical image of the perovskite CsPbBr3 microplate with the twisted moiré potential lattice is shown in Figure 1(b). The parameters of the rhombus are shown in the right panel of Figure 1(b), which induces different periodicities along the horizontal ($x$) and vertical ($y$) directions.
In the experiments, we place the sample such that the moiré pattern is along the vertical ($x$) direction (the long symmetry axis of the rhombus is along the $y$ direction). The energy (E) vs wavevector ($k_{x}$) dispersion of the microcavity is measured by using a home-made angle resolved spectroscopy. The pumping source is a femtosecond laser (5700 Hz) with the wavelength of around 400 nm and pulse width of 130 fs. The pumping spot size of the laser is around 40 $\mu$m which covers around 10 periods of the moiré lattice. The dispersion along $k_{x}$ direction (moiré pattern) and $k_{y}$ direction (perpendicular to the moiré pattern) are shown in Figure 2(a) and Figure 2(b). Firstly, we note that there exist several sets of lower branch polaritons (for simplicity, we only plot LP0 and LP1 in Figure 2(a-b), a larger-range dispersion is plotted in the SM) thanks to the long cavity length where multiple cavity modes strongly couple with the exciton resonance in the perovskite microplates. The strong coupling is confirmed at the large wavevector region where the dispersions are clearly bent when approaching the exciton resonance at around 2.43 eV. Other modes with a larger curvature come from the leakage of photons out of the perovskites or the luminescence from the photoresist.
Figure 2: Exciton polariton dispersions in the moiré lattice. Dispersion along (a) $k_{x}$ direction and (b) $k_{y}$ direction of the moiré lattice. The dash lines indicate the edge of the first Brillouin zones. (c, d) Zoomed dispersion of (a, b) with simulated bands (dash curves) on top.
In addition, in our microcavity, the depth of the moiré potential site is around 20 meV, where we observe both the discrete energy levels (the ground state at 2.26 eV, first excited state at 2.27 eV) and the band structure (annotated in Figure 2(a)) along $k_{x}$ direction for polariton branch LP1. Within this branch, the discrete energy levels are formed due to small overlap of the lower-energy state between moiré potential traps, and polariton bands originate from larger overlap of the higher-energy modes among the moiré potential sites. The discrete polariton energy levels and bands are not visible in LP0 due to negligible photon components.
The size of the Brillouin zone is reversely proportional to the period of the 1D potential lattice. Due to different periods in the real space along $x$ direction and $y$ direction of the moiré lattice, the polariton bands show clear different periodicities and amplitudes in Figure 2(a) and Figure 2(b). The change in the size of the first Brillouin zone (from 1.10 $\mu$m-1 to 1.29 $\mu$m-1) due to different periods along these two directions (from 5.74 $\mu$m to 4.89 $\mu$m) confirms the existence of the coupling between the moiré potential traps along the $x$ direction. Thanks to the coupling strength between the moiré sites, we observe a band gap opening between the ground band and excited band in Figure 2(a) (the ground band structure is not visible due to limited resolution).
We fit the band structure of the moiré lattices and plot the simulated bands in Figure 2(c) which agree with the experimental results very well (the details of fitting is shown in the SM). In addition, the band gap opening is also clearly observed along $k_{y}$ direction, as shown in Figure 2(b). This can be simulated as well, as shown in Figure 2(d). Similar band opening is observed for another lower branch polaritons (shown in SM).
Figure 3: Moiré exciton polariton patterns at different energies. (a-c) Moiré polariton spatial distributions of the ground state, the first excited state and second excite states shown in Figure 2(a). (d-f) Moiré polariton spatial distributions at the bottom of the band 1, the top of the band 1 and the bottom of the band 2 in Figure 2(b). The dash lines represent the two photonic lattices.
Figure 4: Moiré polariton pattern modulation by tuning one photonic lattice. (a,b) Moiré polariton spatial distributions at the ground state and first excited state. The dispersion is shown in the SM.
To check the existence of the moiré patterns, we measure the energy-resolved polariton distribution in the real space. Firstly we show the moiré patterns corresponding to the ground state, the first excited states, which are plotted in Figure 3(a-b). These modes are confined in the single moiré potential traps, where the ground states have a gaussian shape (the emission at the barrier region comes from the bands from LP2, as shown in the SM), the first excited modes are two lobe states with the orientation along $x$ direction. The orientations of the two 1D periodic lattices are indicated by the dash lines. From these graphs, we can observe clear moiré stripes with different patterns along $x$ direction. Due to small overlap between these states trapped in the single moiré sites, the dispersions corresponding to these modes are discrete energy levels. The real space images of the modes with higher energy (2.28 eV) is plotted in Figure 3(c) which corresponds to the ground band along $k_{x}$ direction. This mode consists of three lobes along $x$ direction and promises larger coupling between the nearest potential traps. With higher energy, excited bands are observed along $k_{x}$ direction in Figure 2(a).
On the other hand, the modes at the bottom, the top of the band 1 and the bottom of the band 2 shown in Figure 2(b) are plotted in Figure 3(d, e, f) (band 1 and 2 correspond to the ground and excited band along $k_{y}$ direction. Due to the anisotropic potential distribution along the horizontal and vertical direction, the moiré polariton bands in Figure 2(a) and Figure 2(b) are different). From these, we find clear zero and $\pi$ phase locking of these high-energy modes across the moiré polariton lattice. Note that the in-phase state will appear when the phase difference of polaritons in the moiré lattice is zero, which will give nonzero intensity distribution with the wavevector of $k_{x}=\pm m\pi/a$ (where $m$ is an even integer). On the other hand, the antiphase state or the $\pi$ state can be observed with the wavevector of $k_{x}=\pm m\pi/a$ (where $m$ is an odd integer) yamamoto . At the energy of 2.30 eV which corresponds to the bottom of the band 1 with the wavevector of 0 and ±$\pm 1.1\ \mu$m-1, Figure 3(d) shows a clear moiré pattern along the $y$ direction. In Figure 3(e) and (f) we can observe different moiré patterns at the top of the band 1 and the bottom of band 2 (the energy of these two states are 2.31 eV and 2.32 eV with the wavevector of ±$\pm 0.55\ \mu$m-1 and ±$\pm 1.65\ \mu$m-1, respectively), which is the antiphase state or $\pi$ state. The above moiré polariton patterns are different due to different phase locking. In the zero phase state, the polariton near field distribution is not zero in both the potential trap and barrier along $y$ direction of the moiré lattice due to finite overlapping of the wavefunction, as plotted in Figure 3(d). Whereas antiphase state shown in Figure 3(e) and Figure 3(f) is composed of high energy mode which shows staggered pattern along $y$ direction. More importantly, these states are mainly located at the barrier region of the moiré lattice with the same periodicity as the low-energy modes in Figure 3(a-b).
In our twisted photonics lattices, the energy depth and width of the moiré potential sites can be tuned by modifying the parameters of one photonic lattice. In this case, the moiré polariton can be manipulated and different states can be observed directly. In the experiment, we replace the photoresist lattice onto the perovskite microplate with PMMA layer (the thickness is around 230 nm) and fabricate the photonic lattice (the period is constant with different duty circle of around 40:60) by using laser direct writing technique (the depth is around 50 nm). In this case, the potential depth and width of the PMMA lattice are smaller than the photoresist lattice. Similar amplitude tuning in twisted photonic lattices has been reported to lead the transition from localization to delocalization of photons photonic moire lattice . In our microcavity, we find that the ground state moiré polariton distribution is in the shape of parallelogram. These moiré polaritons are fragile against defects or disorder and become deformed with randomly distribution, for example, dots trapped in the moiré potential, dimers, or delocalized lines, as shown in Figure 4(a). Excited moiré polaritons in Figure 4(b) are robust with disorder in the shape of delocalized lines along the direction of the bottom photonic lattice. These results reveal how the potential depth and width distribution within the twisted photonic structure influence the distribution of the moiré polaritons, and they offers to investigate novel phase transition modulated by the polariton nonlinearity in the future.
Compared with the exciton polaritons due to the strong coupling between the moiré excitons in the TMD monolayer heterostructures and the Fabry-Perot microcavity studied at low temperature moire polariton , we can directly observe the polariton band structure originating from the strong coupling between the excitons and photonic bands which are created by the interaction among the moiré potential traps at room temperature. This allows to tune the polariton moiré bands by manipulating the periodicity, the potential depth and other parameters of the photonic structure much more easily. In addition, although polaritons can be loaded into one dimensional or two dimensional photonic lattices by either etching the pillar or growing buried mesa arrays polariton lieb1 ; polariton lieb2 , our results show that the band structure and spatial modes of the moiré polaritons can be tuned by the particular structure of the twisted potential lattices.
To summarize, we observe the strong coupling between the excitons of the CsPbBr3 microplates and photonic bands of the moiré lattice in a microcavity at room temperature. Different moiré stripes corresponding to different discrete energy levels and polariton bands are observed. The moiré exciton polariton array with different locking phase lays the foundation to study the quantum phase transition between the superfluid and Mott insulators, the Winger crystal or Luttingger liquid in the twisted lattice at room temperature.
Acknowledgments
Acknowledgements.
T.Gao acknowledges the support from the National Natural Science Foundation of China (NSFC, No. $12174285$), especially T.Gao wants to thank Yongyou Zhang and Yang Yang for fruitful discussion. The Paderborn group acknowledges support from the German Research Foundation (DFG, No. 519608013). H.Dai thanks the National Natural Science Foundation of China (NSFC, No. $62375200$).