### v2

# Moiré exciton polaritons in twisted photonic lattices at room temperature

Abstract
Moiré lattices attract significant attention in double-layer graphene and TMD layer heterostructures as well as in photonic crystals due to the interesting exotic physics that emerges within these structures. However, direct measurement of the moiré ground and excited states and Bloch bands in twisted photonic lattices is still illusive. In this work we report strong coupling between excitons in CsPbBr3 microplates and moiré photonic modes at room temperature. Depending on the coupling strength between the nearest potential sites, we observe staggered moiré polariton ground states and excited states trapped in the potential sites and moiré polariton bands across the twisted photonic lattice. Phase locked moiré zero (in-phase) states and moiré $\pi$ (antiphase) states with different spatial distributions are measured. The moiré polariton distribution can be tuned into the shape of a parallelogram by controlling the depth and width of the potential in one photonic lattice with another superimposed one fixed. In addition, moiré polaritons in twisted 2D honeycomb lattices are also observed. Our work lays the foundation to study moiré exciton polariton physics with the potential to observe Wigner crystals and Luttinger liquids in twisted photonic lattices at room temperature.
Moiré lattices in graphene bilayer and TMD monolayer heterostructures offer a tunable platform where the interlayer coupling of electrons and excitons leads to a plethora of physical phenomena, for example, the transition from superconductor to Mott insulator [1], observation of the quantum Hall effect [2] and nontrivial gauge field [3], the appearance of moiré excitons [4, 5, 6], and emergence of correlated exciton insulators [7, 8, 9]. On the other hand, the bands resulting from the coupling between potential sites arranged on a moiré lattice have been extended to photonic systems. In these systems, the localization and delocalization of light and nonlinear optical solitons can be observed by adjusting one photonic lattice against a second one into a particular angle [10]. The period of the moiré lattice in these twisted photonic structures can reach the size of several micros, thus the bands due to the coupling between the moiré potentials can be probed within several $\mu$m-1 in angle-resolved spectroscopy. In this case, both the momentum space and real space imaging can be directly measured. For example, photonic bands in the optical frequency range are probed in a double photonic crystals [11]. Cold atom condensates loaded in twisted photonic lattices have also been realized and the transition from superfluid to Mott insulator in the created moiré lattice is shown [12].
However, direct measurement of the moiré ground state, excited states and moiré band states remain unexplored in the normal twisted photonic lattices. The observation of these modes can be used to investigate distinct state distribution in the momentum and real space resulted from different coupling strength among the moiré potentials. In addition, phase locking between different potential sites can be observed in the moiré bands. For example, the phase synchronization within a moiré flat band can enhance the performance of nanocavity laser array greatly [13]. On the other hand, $\pi$ phase locking at the top of the moiré ground bands and bottom of the moiré excited bands will lead to different mode distribution compared to the ground band, which offers to study quantum simulation with tunable phase modulation. These $\pi$ phase locked moiré states have not been observed yet.
Figure 1: Schematics of the microcavity and optical image of the moiré lattice. (a) The structure of the microcavity. (b) The optical image of the perovskite microplates within the microcavity, the right panel is the zoomed schematics of the moiré pattern.
As a quasiparticle due to the strong coupling between the exciton and photon mode [14], in Fabry-Perot microcavity, exciton polaritons can demonstrate similar Bose-Einstein condensation process as the cold bosonic particles at much higher temperature due to the very light effective mass and large exciton binding energy [15, 16, 17]. More importantly, exciton polaritons allow to investigate different phase locking when they are loaded into a photonic lattice [18]. To observe phase locking of zero and $\pi$ of exciton polaritons in the moiré lattice at room temperature, specific semiconductor materials with robust excitons and easy integration with photonic structures are needed, among which the perovskite CsPbBr3 microplates stand out with large exciton binding energy and oscillator strength. Notably room temperature exciton polariton condensation based on CsPbBr3 microplates has been demonstrated [19, 20, 21, 22, 23, 24].
In the present work we realize and observe strong coupling between the excitons in the perovskite CsPbBr3 microplates and moiré photonic modes in twisted photonic lattice at room temperature. These lattices are composed of two tilted one (two) dimensional lattices fabricated within the microcavity. In our work, the moiré exciton polariton zero state and $\pi$ state in these twisted photonic lattices are measured directly. The staggered moiré polaritons with different energies are either confined in the potential trap or in the potential barrier region. By modifying the potential trap depth and width of one photonic lattice which results in a moiré lattice in the shape of parallelograms, moiré polaritons with the distribution of localized dots, dimers and extended lines are observed due to the sensitivity against disorder. Furthermore, we demonstrate and realize the extension of strong coupling between excitons and moiré photonic bands to twisted 2D honeycomb lattices. Our work gives clear experimental evidence of moiré polaritons at room temperature, and paves the way for future investigations into room temperature quantum simulation and Wigner crystal formation based on moiré exciton polaritons.
The microcavity is fabricated by transferring the chemical vapor deposition (CVD) grown perovskite microplates CsPbBr3 onto the bottom DBR. Firstly, we discuss the interaction of excitons with twisted 1D photonic lattices, where the 1D photoresist lattice with the period of 4 $\mu$m (duty circle: 50:50) and the thickness of 200 nm is formed by using the lithography technique onto the top of perovskite microplates (A layer of SiO2 is deposited onto the perovskite for protection, details can be found in the Methods section). The same 1D photoresist lattice with the same duty circle and depth is formed onto the top DBR with a specific angle against the bottom DBR. The microcavity is fabricated by pasting the top DBR against the bottom DBR using adhesive with an air gap of around 3 $\mu$m. The schematic of the microcavity and the photonic lattice with the twisted angle of around 40 are shown in Figure 1(a), where the moiré stripes consisting of periodic rhombus arrays along horizontal and vertical direction are clearly seen (right panel of Figure 1(a); similar moiré patterns with two identical 1D photonic lattices have been reported in [25]). The optical image of the perovskite CsPbBr3 microplate with the twisted moiré potential lattice is shown in Figure 1(b). The parameters of the rhombus are shown in the right panel of Figure 1(b), which induces different periodicities along the horizontal ($x$) and vertical ($y$) directions.
Figure 2: Photoluminescence exciton polariton dispersions in the moiré lattice. Photoluminescence dispersion along (a) $k_{x}$ direction and (b) $k_{y}$ direction of the moiré lattice. The dash lines indicate the edge of the first Brillouin zones. The inserted graphs in (a, b) are the zoomed dispersions with simulated bands (dash curves) on top.
In the experiments, we place the sample such that the moiré pattern is along the vertical ($x$) direction (the long symmetry axis of the rhombus is along the $y$ direction). The energy (E) vs wavevector ($k_{x}$) dispersion of the microcavity is measured by using a home-made angle resolved spectroscopy. The pumping source is a femtosecond laser (5700 Hz) with the wavelength of around 400 nm and pulse width of 130 fs. The pumping spot size of the laser is around 40 $\mu$m which covers around 8-10 periods of the moiré lattice. The dispersion along $k_{x}$ direction (moiré pattern) and $k_{y}$ direction (perpendicular to the moiré pattern) are shown in Figure 2(a) and Figure 2(b). Firstly, we note that there exist several sets of lower branch polaritons (for simplicity, we only plot LP0 and LP1 in Figure 2(a-b), a larger-energy-range dispersion is plotted in the SM) thanks to the long cavity length where multiple cavity modes strongly couple with the exciton resonance in the perovskite microplates. The strong coupling is confirmed at the large wavevector region where the dispersions are clearly bent when approaching the exciton resonance at around 2.43 eV. Other modes with steeper curvatures come from the leakage of photons out of the perovskites or the luminescence from the photoresist.
Figure 3: Moiré exciton polariton patterns at different energies. (a-c) Moiré polariton spatial distributions of the ground state, the first excited state and ground band states shown in Figure 2(a). (d-f) Moiré polariton spatial distributions at the bottom of band 1, the top of band 1 and the bottom of band 2 in Figure 2(b). The dash lines represent the two photonic lattices.
In addition, in our microcavity, the depth of the moiré potential site is around 20 meV, where we observe both the discrete energy levels (the ground state at 2.26 eV, first excited state at 2.27 eV) and the band structure (annotated in Figure 2(a)) along $k_{x}$ direction for polariton branch LP1. Within this branch, the discrete energy levels are formed due to small overlap of the lower-energy state between moiré potential traps, and polariton bands originate from larger overlap of the higher-energy modes among the moiré potential sites. The discrete polariton energy levels and bands are invisible in LP0 due to much smaller photon components.
The size of the Brillouin zone is reversely proportional to the period of the 1D potential lattice. Due to different periods in the real space along $x$ direction and $y$ direction of the moiré lattice, the polariton bands show cleary different periodicities and amplitudes in Figure 2(a) and Figure 2(b). The change in the size of the first Brillouin zone (from 1.10 $\mu$m-1 to 1.29 $\mu$m-1) due to different periods along these two directions (from 5.74 $\mu$m to 4.89 $\mu$m) confirms the existence of the coupling between the moiré potential traps along the $x$ direction. Thanks to the coupling strength between the moiré sites, we observe a band gap opening between the ground band and excited band in Figure 2(a) (the ground band structure is invisible due to limited resolution).
We fit the band structure of the moiré lattices and plot the simulated bands in the inserted graph of Figure 2(a) which agree very well with the experimental results (the details discussed in SM). In addition, the band gap opening is also clearly observed along $k_{y}$ direction, as shown in Figure 2(b). This can be simulated as well, as shown in the inserted graph of Figure 2(b). Similar band opening is observed for another lower polariton branch (shown in SM).
To check the existence of the moiré patterns, we measure the energy-resolved polariton distribution in real space. Firstly we show the moiré patterns corresponding to the ground state and the first excited states, which are plotted in Figure 3(a-b). In these graphs, we can observe clear moiré states with different patterns along $x$ direction. These modes are confined in the single moiré potential traps, where the ground states have a Gaussian shape (the emission at the barrier region comes from the bands from LP2, as shown in the SM), and the first excited modes are two-lobe states with the orientation along $x$ direction. The orientations of the two 1D periodic lattices are indicated by the dashed lines. Due to small overlap between these states trapped in the single moiré sites, the dispersions corresponding to these modes are discrete energy levels. The real space image of the modes with a larger energy (2.28 eV) is plotted in Figure 3(c) which corresponds to the moiré ground band along $k_{x}$ direction. This mode consists of three lobes along $x$ direction and promises larger coupling between the nearest potential traps. With higher energy, moiré excited bands are observed along $k_{x}$ direction in Figure 2(a).
On the other hand, the moiré polariton modes at the bottom (top) state of the band 1 and the bottom state of the band 2 shown in Figure 2(b) are plotted in Figure 3(d, e, f) (band 1 and 2 correspond to the ground and excited band along $k_{y}$ direction. The moiré polariton bands in Figure 2(b) are different from Figure 2(a) due to the anisotropic potential distribution along the horizontal and vertical direction of the twisted photonic lattices). From these, we find clear zero and $\pi$ phase locking of these high-energy modes across the moiré polariton lattice. Note that the in-phase state will appear when the phase difference of polaritons in the moiré lattice is zero, which will give nonzero intensity distribution with the wavevector of $k_{y}=\pm m\pi/a$ (where $m$ is an even integer). On the other hand, the antiphase state or $\pi$ state can be observed with the wavevector of $k_{y}=\pm n\pi/a$ (where $n$ is an odd integer) [18]. At the energy of 2.30 eV which corresponds to the zero phase state located at the bottom of band 1 with the wavevector of 0 and ±$\pm 1.1\ \mu$m-1, Figure 3(d) shows a clear moiré pattern along the $y$ direction. In Figure 3(e) and (f) we can observe different moiré patterns at the top of the band 1 and the bottom of band 2 (the energy of these two states are 2.31 eV and 2.32 eV with the wavevector of ±$\pm 0.55\ \mu$m-1 and ±$\pm 1.65\ \mu$m-1), which is the antiphase state or $\pi$ state. In above moiré zero phase state, the polariton near field distribution is visible in both the potential trap and barrier along $y$ direction of the moiré lattice due to finite overlap of the wavefunctions, as plotted in Figure 3(d). Whereas the moiré antiphase states shown in Figure 3(e) and Figure 3(f) are composed of a higher energy mode which shows a staggered pattern along $y$ direction. More importantly, these states are mainly located at the barrier region of the moiré lattice with the same periodicity as the low-energy modes in Figure 3(a-b).
The moiré polaritons can be engineered quite easily within our photonic structures. In the twisted 1D lattices, the energy depth and width of the moiré potential sites can be tuned by modifying the parameters of one photonic lattice. In this case, different moiré polariton states can be observed directly. In the experiment, we replace the photoresist lattice on the perovskite microplate with a PMMA layer (the thickness is around 230 nm) and fabricate the photonic lattice (the period is still 4 $\mu$m, but the potential trap width is 1.6 $\mu$m, the potential barrier with is 2.4 $\mu$m) by using laser direct writing technique (the depth is around 50 nm). In this case, the potential depth and width of the PMMA lattice are smaller than the photoresist lattice. Similar amplitude tuning in twisted photonic lattices has been reported to lead to a transition from localization to delocalization of photons [10]. In our microcavity, we find that the moiré polariton ground state distribution is in the shape of a parallelogram. These moiré polaritons are fragile against defects or disorder and become deformed with random distribution, for example, localization trapped in the moiré potential, dimers, or delocalized lines, as shown in Figure 4(a). Excited moiré polariton modes in Figure 4(b) are robust with disorder in the shape of delocalized lines along the direction of the top photonic lattice. Finally, our experimental scheme to observe moiré polaritons can be extended to twisted 2D honeycomb lattices. To realize this, we fabricate a 2D honeycomb lattice onto the top of the perovskite micrplates and the top DBR using a similar method. With the twisting angle of 11 as shown in Figure 5(a), we can observe different moiré polariton patterns as shown in Figure 5(b-d) (the dispersions are shown in the SM). The observation of these tunable moiré polaritons enables studies of novel polariton phase transitions at room temperature in future works.
Figure 4: Moiré polariton patterns modulation by tuning one photonic lattice. (a) Schematics of the moiré pattern in two different twisted photonic lattice. The right panel indicates the specific parameter of the supercell. (b, c) The spatial distribution of the moiré polariton ground state and excited state, where the dispersion of the microcavity is shown in SM.
Figure 5: Moiré polariton patterns in twisted 2D photonic lattice. (a) Schematics of the moiré pattern in twisted 2D lattices. (b-d) The spatial distribution of the moiré polariton ground state, first excited state and second excited state, with the dispersion of the microcavity shown in SM.
Compared with the exciton polaritons due to the strong coupling between moiré excitons in TMD monolayer heterostructures and Fabry-Perot microcavity studied at low temperature [26], we can directly measure the moiré polariton ground state, moiré excited state and moiré polariton bands in twisted photonic lattices at room temperature. This allows to observe how the moiré polaritons are manipulated by the periodicity, the potential depth and other parameters of the photonic structure much more easily. In addition, although polaritons can be loaded into 1D or 2D photonic lattices by either etching pillars or growing buried mesa arrays [27, 28], our results show that the bands and spatial modes of the moiré polaritons can be observed and engineered by specific structures of twisted potential lattices, this provides to study moiré polariton based integrated photonic chips and circuits.
To summarize, we observe strong coupling between excitons of CsPbBr3 microplates and moiré photonic bands of 1D (2D) twisted lattice in a microcavity at room temperature. Different moiré polariton stripes corresponding to the discrete energy levels and the continuous bands are observed. The realization of moiré exciton polariton arrays with different phase locking can be used to investigate quantum phase transitions between the superfluid and Mott insulators, Winger crystals and Luttinger liquids in twisted photonic lattices at room temperature.
Acknowledgements.
T.Gao acknowledges the support from the National Natural Science Foundation of China (NSFC, No. $12174285$), especially T.Gao wants to thank Yongyou Zhang and Yang Yang for fruitful discussion. The Paderborn group acknowledges support from the German Research Foundation (DFG, No. 519608013). H.Dai thanks the National Natural Science Foundation of China (NSFC, No. $62375200$).